OpenStudy (asylum15):

It takes a horizontal force of 375N to drag a box of mass 300kg at uniform speed along a level track. Calculate the coefficient of friction between the contact surfaces.

OpenStudy (asylum15):

mu = f/n?

OpenStudy (anonymous):

Friction force = N * mu N = mg Friction force = - 375 N

OpenStudy (asylum15):

Archie didn't read that properly, lol. Looking for mu :)

OpenStudy (anonymous):

But is that so difficult to derive mu from given formulas?

OpenStudy (anonymous):

))

OpenStudy (asylum15):

See my first message? I was wondering if I was correct ;) If the box was then inclined at 40 degrees, the magnitude of the force parallel to the slope would be w.sin.theta?

OpenStudy (anonymous):

F = ma = 0. F - f = 0, so F = f, where f is frictional force. f = uN. 375N = u(300kg)(9.8m/s^2). u = .13

OpenStudy (asylum15):

Any ideas for the 2nd part Aj? :)

OpenStudy (anonymous):

\[\mu \]=375\[\div \]3000 hence \[\mu \] = 0.125